OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region bounded by the given curves about the specified axis?

OpenStudy (anonymous):

\[y=\frac{ 1 }{ x^3 }, y=0, x=3, x=6\]about the y-axis

OpenStudy (anonymous):

I haven't done this in a while, but the point of intersection is (3,1) so it should be: pi*Integral( [3y^2]^2,y,0,1)=1.8*pi=5.654 since the rule is pi*integral( upper^2-lower^2,variable,lower x or y bound, upper x or y bound)

OpenStudy (anonymous):

Shouldn't the integration be from 3 to 6?

OpenStudy (campbell_st):

no... you substiute x = 3 and x = 6 to find the corresponding y values so you are integrating from y = 1/216 to 1/27 you will also need to make x the subject... then it \[V = \pi \int\limits_{\frac{1}{216}}^{\frac{1}{27}} x^2 dy\]

OpenStudy (campbell_st):

the diagram is